NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 19:02:22 on September 14, 2009
Entry number 289546
keyword=Run Coordinator's Report for September 13 Swing to September 14 Day
Run Coordinator's Report for September 13 Swing to September 14 Day
Summary
Production running and Compton data-taking throughout the night. We had
an access to reboot the Compton HV crate, and an access to fix the left
Q1. We are still having some issues with the Compton background rates,
but we are taking data with it, and the experts are working on
understanding these data.
Significant News
- Production running and Compton data-taking last night. We
were limited to 80 muA for an hour while MCC did tests to determine the
problem. When beam delivery resumed, it was more stable and we went back
to 100 muA. We lost a couple of hours on owl shift due to a problem at
the MCC, but the beam came back at about 8:00am looking pretty good.
- We took an access at 5:10am because the Compton photon HV crate had
locked up and needed to be rebooted. We took a quick access and
Alexandre rebooted it.
- The MCC took the beam for tuning at 9:45am and finished at 10:19am,
but the tune was not good for the Compton, so the beam was retuned. At
11:40am we went back to the original tune.
- At 2:00pm, the Compton rates crept up to the point that the detector
HV was tripping off. The MCC operator said that he saw that our Compton
is having issues, but nothing had been changed on his end. They worked
on adjusting it, talking to Alexandre so that they can communicate with
an expert.
- Also at 2:00pm, the left Q1 started behaving bizarrely, with the
momentum readback fluctuating around, but the other readbacks not
changing, and with the alarm for it going off. We requested a controlled
access, and Jack and Heidi repaired it (halog
entry 289528 and halog
entry 289529).
- The Compton is still having issues, but we are now taking Compton
data, and the experts are working on analyzing and understanding these
data so that we have a better understanding of what is going on.
Plans
- Take lots of production data! Watch the lumi widths to see if
anything changes. Please check the histograms on the online data monitor
with the reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
- Take Compton data.
- The beam will be taken away for the C-50 commissioning at 7:00am.
The hall will go into controlled access after that. Please contact me if
you want to do work in the hall and haven't already let me know.
Concerns
- Please write a detailed shift summary (including runs taken and any
activities) and submit the BTA for each shift.
- Please make sure that the shift checklist is done once each shift.
- The Compton polarimeter (but they are taking data now).
- There are lots of empty shifts that need to be filled immediately!
Please sign up!
- Keep a close eye on the lumi widths as we take data.
- A WAC schedule is being created and will be posted so that I know
who is doing this task.
- Beam Time Accounting: take off 1 minute per beam trip, since the BTA
code doesn't.
- The dithering looks non-invertible. Someone should look at this.
- There is a constant alarm due to unused VDCs. This alarm should be
removed so that people don't become desensitized to the alarm handler.
Coulombs So Far
- Coulombs taken in this time period: 5.1 C
- Coulombs taken in all so far: 32.9 C as of 14:44
Shift Summaries and BTA Reports